OpenStudy (anonymous):

A water cooler holds 25 liters of sports drink. Approximately how many gallons is this? (1 gallon = 3.785 liters) 5 gallons 6 gallons 7 gallons 21 gallons

OpenStudy (anonymous):

@MissSmartiez

OpenStudy (fortytherapper):

Let's cross multiply, then divide When doing so, you wanna keep it consistent \[\frac{ 1 }{ 3.785 } = \frac{ x }{ 25 }\]

OpenStudy (xmissalycatx):

I've literally already helped you...

OpenStudy (xmissalycatx):

@FortyTheRapper I've already taught her that, the only thing she is having trouble with is dividing the 25 and 3.785

OpenStudy (fortytherapper):

Calculators are handy also, then

OpenStudy (xmissalycatx):

She was using one, but she kept dividing it wrong.

OpenStudy (anonymous):

=(

OpenStudy (fortytherapper):

Hmm Let's form the cross in "cross" multipying. Make a cross and connect the dots to get \[(25)(1) = (3.785)(x)\] Now if we divide both sides by 3.785, we can cancel out what's on the right because those are the same number, yes? If we also divided the left by 3.785, the top number is he first number you put into the calculator

OpenStudy (anonymous):

May you tell me the answer i prmise this is literly the first time you will do it i just need to get throu this please.

OpenStudy (fortytherapper):

Just put \[\frac{ 25 }{ 3.785 } \] into the calculator and see what you get

OpenStudy (damonte212):

..